OpenStudy (anonymous):

A 9500 kg boxcar traveling at 17 m/s strikes a second boxcar at rest. The two stick together and move off with a speed of 7.0 m/s. What is the mass of the second car?

OpenStudy (anonymous):

let m1 be the mass of first car and let m2 be the mass of the second car.From the law of conservation of momentum,momentum before=momentum afte

OpenStudy (anonymous):

therefore,\[m1\times v +m2 \times its velocity\]

OpenStudy (anonymous):

will be equal to total mass times velocity

OpenStudy (anonymous):

therefore,\[9500\times17 + m2\times 0=(9500+m2)\times 7\]

OpenStudy (anonymous):

solve that and u get ur answer